Waller, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 15 years to maturity that is quoted at 107 percent of face value. The issue makes semiannual payments and has an embedded cost of 7 percent annually. The company's pretax cost of debt is _______ percent. If the tax rate is 35 percent, the aftertax cost of debt is ________ percent.
Answer #1
According to the given information,
Number of years to maturity = 15
Face value = $1000
Current value = $1070
Coupon rate = 7%
Semi-annual coupon rate = 3.5%
Semi-annual coupon payment = 3.5% * $1000
                                             = $35
Calculating the pre-tax cost of debt using excel sheet:
Step1: Go to excel and click "insert" to insert the function.
Step2: Select the "Rate" function as we are finding the pre-tax cost of debt in this case.
Step3: Enter the values as Nper = 15 * 2 ; PMT = -35; PV = 1070; FV = -1000
Step4: Click "OK" to get the desired value.
The value comes to "3.14%"
Annual pre-tax cost of debt is 6.28%
After-tax cost of debt = Pre-tax cost of debt (1 - Tax rate)
                                 = 0.0628 (1 - 0.35)
                                 = 0.0628 (0.65)
                                 = 0.04082 or 4.082%
Therefore, the after-tax cost of debt is 4.082%